# Coward: Toward Practical Proactive Federated Backdoor Defense via Collision-based Watermark

### v1

@@ V-E Why is Our Method Effective?
Case Study. To intuitively explain how our method mitigates OOD bias, we present a set of representative detection cases of our method in Figure 8, ranging from an ideal non-bias scenario to increasingly biased conditions. Each case visualizes the prediction confusion matrices of both raw and watermarked OOD samples, revealing how local training and OOD bias impact the two injected mappings. Specifically, the top two rows show the prediction results of raw OOD samples, while the bottom two rows show those of watermarked samples. Each case includes a small subfigure illustrating the cumulative distribution across target classes. To aid understanding, we highlight that the attack target is class 0, and the watermark target is class 8. We elaborate details as follows:
(1) Ideal Cases: Case 1 illustrates a malicious client (C0), where both the OOD and watermark mappings collapse into the attack target class 0, indicating severe OOD mapping distortion and a strong collision effect caused by the backdoor. This results in low watermark accuracy, enabling effective detection. In contrast, Case 2 represents an ideal benign client unaffected by OOD bias, where both the OOD (diagonal) and watermark mappings are fully preserved. Empirically, OOD bias seldom overrides backdoors but often impacts benign clients; Thus we examine biased cases for benign clients below. (2) Biased Cases: Case 3 presents a mild OOD bias where the OOD mapping is largely retained but partially biased toward class 9, yet the watermark prediction remains dominated by class 8. This highlights the strength of the watermark, which consistently guides triggered OOD samples to the target class, even when raw samples are misclassified. In Case 4, under a stronger bias where OOD predictions collapse into two dominant classes, the watermark still holds the highest confidence for class 8, indicating resilience even under severe distributional shifts. Intuitively, after watermark planting, the model forms a direct association between the trigger pattern and the target label, making it less influenced by the background image and thus more robust to OOD bias. Case 5 presents an extreme scenario where the OOD mapping collapses to a single class that coincidentally aligns with the watermark target. These cases collectively indicate our robustness against different degrees of OOD bias.
@@
VI Conclusion
In this paper, we revisited backdoor defenses in federated learning, revealing the non-i.i.d. blurring and out-of-distribution (OOD) prediction bias that had constrained prior passive and proactive approaches. To tackle these challenges, we designed a new method, Coward, which leverages a collision-based OOD watermark that the server injects and later inspects to expose malicious clients, inspired by the phenomenon of multi-backdoor collision effects. By treating high, rather than low, watermark accuracy as the sign of benign behavior, Coward mitigated the side effects of OOD bias. Extensive experiments verified the effectiveness of Coward and its resilience against potential adaptive attacks. Our method provides a new perspective for understanding and defending backdoor attacks in federated learning, facilitating more secure and trustworthy data sharing and decentralized training.
@@ Appendix A Collision Effects on OOD Watermark
In Section III-C, we introduced the multi-backdoor collision effect under a centralized setting where two backdoors share the same dataset, aiming to illustrate its core mechanism. To further demonstrate how this effect manifests in our OOD-based watermark under practical FL scenarios, we provide an extended analysis in this section. Specifically, we examine the collision effect between the attacker’s backdoor and our OOD watermark across progressively complex scenarios: starting from a centralized setup, moving to a static FL environment with fixed client participation, and finally to the dynamic FL scenario adopted in our main experiments. To align with main experiments, we conduct all evaluations under the 0.9 non-i.i.d. setting on the CIFAR-10 dataset by default. As elaborated in the following sections, results across all three stages consistently reveal a clear collision effect, validating the capability of our OOD watermark to trigger such interactions and effectively distinguish malicious clients.
@@
Centralized Scenario. We conduct a similar experiment as in Section III-C, with the key difference that the second backdoor injection is replaced by our OOD watermark planting process. Specifically, we use a target label 8 watermark with MNIST as the planting set. As shown in Fig. 2, in each subfigure, the left side illustrates the collision effect, while the right side shows the discrepancy between benign ($i.e.$, finetuning) and malicious ($i.e.$, backdoor injection) operation. Specifically, in subfigure (a), we observe a clear collision effect on the OOD watermark, characterized by a clear drop in ASR and a notable distinction between benign and malicious clients. In subfigure (b), when the BN layer switch is disabled, the benign-malicious discrepancy becomes indistinguishable, indicating the necessity of BN isolation for effective detection.
@@
Static FL Scenario. We consider a slightly more complex scenario under a static FL setting, involving 10 clients with full participation in each communication round. Among them, only one client is malicious. This simplified full-participation setup facilitates the observation of each client’s temporal behavior and helps illustrate the collision effect in a federated learning context. In each global round, the server injects an OOD watermark, while all benign clients perform standard local training.To clearly observe the collision effect, the attacker is configured to behave benignly for the first 20 rounds and initiates a backdoor attack targeting label 0 from round 21 onward. After each local update, we inspect the watermark accuracy of all 10 participated clients. As shown in Figure 3(b), once the malicious client begins the attack, its watermark accuracy drops sharply to nearly zero, indicating a strong collision effect. In contrast, the remaining nine benign clients consistently maintain high watermark accuracy. This sharp divergence highlights the benign-malicious discrepancy induced by the collision effect, which is critical for enabling effective backdoor detection. Additionally, in subfigure (a), we observe that benign clients struggle to stably preserve the watermark without BN layer isolation, underscoring the importance of BN switching in reducing FPR.
@@
Dynamic FL Scenario. We finally present the collision effect under the dynamic FL setting used in our main experiments, where 10 out of 100 clients are randomly selected to participate in each global round. This dynamic participation leads to divergent and inconsistent watermark retention behaviors among benign clients, making direct comparisons between them infeasible. Therefore, we focus on observing relative trends across rounds. Specifically, we plot the temporal variation of watermark accuracy for all clients based on their actual participation order. To control page length, we select clients at intervals of 3 ($i.e.$, client IDs with a gap of 3). As shown in Figure 1, we observe the following: (1) Malicious client 0 exhibits a pronounced collision effect, with watermark accuracy dropping to nearly zero. (2) Benign clients maintain relatively high watermark accuracy (mostly above 50%), though the magnitude and variation patterns differ across clients due to heterogeneous participation and client distribution. Overall, the results reveal a clear separation between benign and malicious clients, enabling robust and tolerant threshold selection. Despite being influenced by dynamic participation and non-i.i.d.data, the collision effect remains highly effective in distinguishing malicious behavior.
@@ Appendix B Discussion on Watermark Configuration
TABLE IV: Performance of Coward under different levels of global model regularization.
$\lambda$ TPR FPR ASR BA
@@ Appendix C Ablation Study on Watermark Injection
Impact of Global Model Regularization. Table IV reports the effect of varying levels of global model regularization. We observe a positive impact on TPR, where applying regularization consistently yields higher TPRs compared to the no-regularization baseline. Increasing the regularization strength $\lambda$ leads to slight improvements in TPR, with the best performance observed at $\lambda=0.3$. These results suggest that a moderate level of regularization is beneficial for achieving optimal detection performance.
Learning Efficiency of Watermark Injection. We conduct ablation studies on three key hyperparameters that control the strength of watermark injection: the number of OOD watermark samples, the watermark learning rate, and the number of injection iterations per global round. We vary one factor at a time while keeping the others fixed. By default, the watermark learning rate is set to 0.001, the number of watermark injection rounds to 5, and the watermark dataset size to 1000. As shown in Table III, Coward consistently achieves strong performance, with TPR exceeding 98% and FPR remaining below 10% in most cases. Even under extremely constrained injection settings ($e.g.$, a watermark dataset size of 200, a learning rate of 0.0001, or a single injection round), the method still maintains competitive results, achieving 100% TPR with FPR below 25%. Overall, effective watermark injection requires only a low level of effort. Most parameter variations lead to only marginal performance differences, and any configuration beyond the most extreme cases yields satisfactory performance. These results indicate that our method is robust across a wide range of watermark injection strengths and introduces negligible computational overhead in federated learning. Notably, effective defense can be achieved with as few as only 1 round of watermark injection.
@@ Appendix D Potential Limitations and Future Directions
[…]

### v2

@@ 5.5 Why is Our Method Effective?
Case Study. To intuitively explain how our method mitigates OOD bias, we present a set of representative detection cases of our method in Figure 8, ranging from an ideal non-bias scenario to increasingly biased conditions. Each case visualizes the prediction confusion matrices of both raw and watermarked OOD samples, revealing how local training and OOD bias impact the two injected mappings. Specifically, the top two rows show the prediction results of raw OOD samples, while the bottom two rows show those of watermarked samples. Each case includes a small subfigure illustrating the cumulative distribution across target classes. To aid intuitive understanding, we highlight that the attack target is class 0, and the watermark target is class 8. We elaborate on the case study by explaining both the ideal unbiased scenario and the biased scenario.:
  • Ideal Cases: Case 1 illustrates a malicious client (C0), where both the OOD and watermark mappings collapse into the attack target class 0, indicating severe OOD mapping distortion and a strong collision effect caused by the backdoor. This results in low watermark accuracy, enabling effective detection. In contrast, Case 2 represents an ideal benign client unaffected by OOD bias, where both the OOD (diagonal) and watermark mappings are fully preserved. Empirically, OOD bias seldom overrides backdoors but often impacts benign clients; Thus we examine biased cases for benign clients below.
  • Biased Cases: Case 3 presents a mild OOD bias where the OOD mapping is largely retained but partially biased toward class 9, yet the watermark prediction remains dominated by class 8. This highlights the strength of the watermark, which consistently guides triggered OOD samples to the target class, even when raw samples are misclassified. In Case 4, under a stronger bias where OOD predictions collapse into two dominant classes, the watermark still holds the highest confidence for class 8, indicating resilience even under severe distributional shifts. Intuitively, after watermark planting, the model forms a direct association between the trigger pattern and the target label, making it less influenced by the background image and thus more robust to OOD bias. Case 5 presents an extreme scenario where the OOD mapping collapses to a single class that coincidentally aligns with the watermark target. These cases collectively indicate our robustness against different degrees of OOD bias.
@@
6 Conclusion
In this paper, we investigated the detection of backdoor manipulation that hinder the trustworthy development of modern AI systems. Focusing on the emerging privacy-preserving federated learning paradigm, we revisited existing FL backdoor detection methods and identified two fundamental limitations: non-IID blurring and out-of-distribution (OOD) prediction bias. To tackle these challenges, we designed a new method, Coward, which leverages a collision-based OOD watermark that the server injects and later inspects to expose malicious clients. Inspired by the multi-backdoor collision effects, we carefully design Coward to enable a transition of the detection signal from low to high watermark accuracy, thereby mitigating the adverse effects of OOD bias. Extensive experiments verified the effectiveness of Coward and its resilience against potential adaptive attacks.
Figure 1: OOD watermark collision under dynamic FL scenario. The collision effect remains highly effective in distinguishing malicious behavior under dynamic federated participation. The attacker exhibits a strong collision effect, while benign clients show diverse but generally higher levels of watermark retention.
(a) w./ BN swith
(b) w./o. BN switch
Figure 2: OOD watermark collision under centralized scenario. Our OOD watermark is planted as the second backdoor. The resulting collision effect is significant, regardless of whether the BN layer is switched. However, switching the BN layer creates a more pronounced performance discrepancy between benign finetuning and backdoor injection.
(a) w./o. BN switch
(b) w. BN switch
Figure 3: OOD watermark collision under static FL scenario. Once the attacker (in red) begins injecting the backdoor at round 1021, its watermark accuracy rapidly drops to 0, indicating a strong collision effect and a clear distinction from benign clients. Moreover, the BN switch plays a critical role in accurately reflecting the benign clients’ (in varying shades of blue) ability to retain the watermark.
(a) attacker triggers
(b) watermark triggers
Figure 4: Visualization of trigger configurations.
@@ Appendix A Collision Effects on OOD Watermark
In Section 3.4, we introduced the multi-backdoor collision effect under a centralized setting where two backdoors share the same dataset, aiming to illustrate its core mechanism. To further demonstrate how this effect manifests in our OOD-based watermark under practical FL scenarios, we provide an extended analysis in this section. Specifically, we examine the collision effect between the attacker’s backdoor and our OOD watermark across progressively complex scenarios: starting from a centralized setup, moving to a static FL environment with fixed client participation, and finally to the dynamic FL scenario adopted in our main experiments. To align with main experiments, we conduct all evaluations under the 0.9 non-i.i.d. setting on the CIFAR-10 dataset by default. As elaborated in the following sections, results across all three stages consistently reveal a clear collision effect, validating the capability of our OOD watermark to trigger such interactions and effectively distinguish malicious clients.
@@
Centralized Scenario. We conduct a similar experiment as in Section 3.4, with the key difference that the second backdoor injection is replaced by our OOD watermark planting process. Specifically, we use a target label 8 watermark with MNIST as the planting set. As shown in Fig. 2, in each subfigure, the left side illustrates the collision effect, while the right side shows the discrepancy between benign ($i.e.$, finetuning) and malicious ($i.e.$, backdoor injection) operation. Specifically, in subfigure (a), we observe a clear collision effect on the OOD watermark, characterized by a clear drop in ASR and a notable distinction between benign and malicious clients. In subfigure (b), when the BN layer switch is disabled, the benign-malicious discrepancy becomes indistinguishable, indicating the necessity of BN isolation for effective detection.
@@
Static FL Scenario. We consider a slightly more complex scenario under a static FL setting, involving 10 clients with full participation in each communication round. Among them, only one client is malicious. This simplified full-participation setup facilitates the observation of each client’s temporal behavior and helps illustrate the collision effect in a federated learning context. In each global round, the server injects an OOD watermark, while all benign clients perform standard local training.To clearly observe the collision effect, the attacker is configured to behave benignly for the first 20 rounds and initiates a backdoor attack targeting label 0 from round 21 onward. After each local update, we inspect the watermark accuracy of all 10 participated clients. As shown in Figure 3(b), once the malicious client begins the attack, its watermark accuracy drops sharply to nearly zero, indicating a strong collision effect. In contrast, the remaining nine benign clients consistently maintain high watermark accuracy. This sharp divergence highlights the benign-malicious discrepancy induced by the collision effect, which is critical for enabling effective backdoor detection. Additionally, in subfigure (a), we observe that benign clients struggle to preserve the watermark without BN layer isolation, underscoring the importance of BN switching in reducing FPR.
@@
Dynamic FL Scenario. We finally present the collision effect under the dynamic FL setting used in our main experiments, where 10 out of 100 clients are randomly selected to participate in each global round. This dynamic participation leads to divergent and inconsistent watermark retention behaviors among benign clients, making direct comparisons between them infeasible. Therefore, we focus on observing relative trends across rounds. Specifically, we plot the temporal variation of watermark accuracy for all clients based on their actual participation order. To control page length, we select clients at intervals of 3 ($i.e.$, client IDs with a gap of 3). As shown in Figure 1, we observe the following: (1) Malicious client 0 exhibits a pronounced collision effect, with watermark accuracy dropping to nearly zero. (2) Benign clients maintain relatively high watermark accuracy (mostly above 50%), though the magnitude and variation patterns differ across clients due to heterogeneous participation and client distribution. Overall, the results reveal a clear separation between benign and malicious clients, enabling robust and tolerant threshold selection. Despite being influenced by dynamic participation and non-i.i.d.data, the collision effect is still effective in distinguishing attackers.
@@ Appendix B Discussion on Watermark Configuration
TABLE VII: Performance of Coward under different levels of global model regularization.
[…]
@@ Appendix C Ablation Study on Watermark Injection
Impact of Global Model Regularization. Table VII reports the effect of varying levels of global model regularization. We observe a positive impact on TPR, where applying regularization consistently yields higher TPRs compared to the no-regularization baseline. Increasing the regularization strength $\lambda$ leads to slight improvements in TPR, with the best performance observed at $\lambda=0.3$. These results suggest that a moderate level of regularization is beneficial for achieving optimal detection performance.
Learning Efficiency of Watermark Injection. We conduct ablation studies on three key hyperparameters that control the strength of watermark injection: the number of OOD watermark samples, the watermark learning rate, and the number of injection iterations per global round. We vary one factor at a time while keeping the others fixed. By default, the watermark learning rate is set to 0.001, the number of watermark injection rounds to 5, and the watermark dataset size to 1000. As shown in Table VI, Coward consistently achieves strong performance, with TPR exceeding 98% and FPR remaining below 10% in most cases. Even under extremely constrained injection settings ($e.g.$, a watermark dataset size of 200, a learning rate of 0.0001, or a single injection round), the method still maintains competitive results, achieving 100% TPR with FPR below 25%. Overall, effective watermark injection requires only a low level of effort. Most parameter variations lead to only marginal performance differences, and any configuration beyond the most extreme cases yields satisfactory performance. These results indicate that our method is robust across a wide range of watermark injection strengths and introduces negligible computational overhead in federated learning. Notably, effective defense can be achieved with as few as only 1 round of watermark injection.
@@ Appendix D Potential Limitations and Future Directions
  • Fine-grained OOD Watermark Design. Our current OOD watermark adopts a simple yet effective configuration applied uniformly across the entire planting set, using widely adopted trigger patterns. Exploring more complex and fine-grained designs that potentially leverage the diversity of OOD samples($e.g.$, sample-specific or class-conditional watermark mappings) would be a valuable future direction.